OpenStudy (anonymous):

to apply herbicide to a large field she must mix a solution of 67% herbicide with a solution of 46% herbicide to form 42 liters of a 55%solution.how much the 67%solution must they use ? A.23L B.18L C.34L D35L

OpenStudy (anonymous):

B

OpenStudy (anonymous):

can you tell me the equation of how to solve it

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

I used this web side to solve eq. http://math.cowpi.com/systemsolver/2x2.html

OpenStudy (mertsj):

x= liters of 67% solution 42-x= liters of 46% solution .67x+.46(42-x)=.55(42) Multiply both sides by 100 to get rid of decimal points: 67x+46(42-x)=55(42) 67x+1932-46x=2310 21x=378 x=18 liters of 67% solution 42-x=24 liters of 46% solution